# NuSTAR Detection of an Absorption Feature in ESP 39607: Evidence for an Ultra-fast Inflow?

### v1

@@ Abstract
We report the serendipitous discovery of an absorption feature at 4.8 keV in the NuSTAR spectra of ESP 39607, a Seyfert 2 galaxy at , observed in May 2023 and August 2024. The feature is detected in both observations with individual significance levels between 2 and 3, computed with multiple statistical methods. The combined probability of detecting it in both observations is 4. The absorption feature is consistent with an ultra-fast inflow (UFI) potentially associated with Fe xxv or Fe xxvi K transitions. The inferred inflow velocity is 0.15–0.20, with an estimated launching radius of 22–89 , depending on the assumed iron transition and whether radiation pressure is accounted for. Photoionization modeling associates the UFI primarily with Fe xxv K absorption, blended with a minor contribution from Fe xxvi K. Alternative explanations, including associations with the warm-hot intergalactic medium or outflows of lighter elements, were investigated but found unlikely. If confirmed, this detection represents a rare example of a UFI, providing valuable evidence into extreme and/or non-standard accretion processes near supermassive black holes. Follow-up observations with higher-resolution X-ray spectroscopy, such as with XMM-Newton or XRISM, will be essential to confirm the nature of this feature and better constrain the physical mechanisms driving it.
@@ 2 Data
In addition to the NuSTAR data analyzed in this work, a 60.7 ks Suzaku observation from December 2010 is also available for ESP 39607. We used the spectra extracted by Ricci et al. (2017) and refer the reader to that work for details on data reduction. The same dataset is also included and shown in the broad-band spectral fitting analysis presented in Peca et al. (submitted), and we refer to both studies for additional modeling details. Applying the same procedure described in Section 3, we did not find evidence of the absorption feature seen in the NuSTAR spectra within the Suzaku data. The implications of this observation are discussed in Section 4.
@@ 3.1 May 2023 observation
We started by fitting the 2023 observation, which has a net (i.e., background subtracted) count rate of 0.0500.002 . First, we modeled the source spectrum using a simple model composed of an absorbed power-law with an energy cut-off fixed at 200 keV (e.g., Ricci et al., 2017; Zappacosta et al., 2018) plus a pexrav (Magdziarz & Zdziarski, 1995) component to account for reflected emission. For simplicity, we assumed pure reflection by fixing (e.g., Piconcelli et al., 2011; Zhao et al., 2019; Serafinelli et al., 2024); the negative value directs the model to operate in reflection-only mode (e.g., Alexander et al., 2011; Boorman et al., 2025), while a value of one corresponds to reflection from an infinite slab illuminated by isotropic coronal emission (e.g., Marchesi et al., 2016; Zappacosta et al., 2018). The inclination angle was set to the default value of 60∘ (e.g., Ueda et al., 2007; Zhao et al., 2020). After verifying consistency between the normalizations of the primary power-law and reflection components, we linked them to the same value to minimize degeneracies (e.g., Liu et al., 2017; Peca et al., 2023b). We refer to this configuration as Model 1. Second, we fitted the source spectrum with the more complex, but physically motivated and self-consistent model, UXCLUMPY (Buchner et al., 2019). We fixed the TORsigma and CTKcover parameters to their standard values, 28 and 0.4, respectively, and the inclination angle to 60∘, as for the first model. We refer to this configuration as Model 2. Results from Model 2 are consistent with those from Model 1 within the uncertainties (see Table 1), and both models confirm the obscured and luminous nature of the source, with derived cm-2 and erg s-1, in agreement with those previously reported by Ricci et al. (2017).
The fit residuals revealed a possible absorption feature at 4.8 keV (Figure 1, left panel). To model this feature, we introduced a Gaussian line component to both models, where line normalization and energy were left free to vary. The line width was fixed to eV (e.g., Tombesi et al., 2010; Vignali et al., 2015), as it was found to be unconstrained at the lower end when left free, i.e., consistent with a narrow and unresolved feature. This inclusion improved the fit significantly. For Model 1, we obtained /2, which corresponds roughly to a fit improvement of 3 assuming Gaussian approximation (e.g., Tozzi et al., 2006). For Model 2, the improvement was /2, i.e., 2.7. Model 1 yielded keV and an equivalent width of keV, with Model 2 giving consistent results. We summarize all the detailed results for all model configurations in Table 1.
@@
To further assess the reliability of the absorption feature detection, we adopted a Monte Carlo (MC) approach, as widely used in the literature to evaluate the significance of emission and absorption lines (e.g., Markowitz et al., 2006; Vignali et al., 2015; Costanzo et al., 2022; Peca et al., 2023a). We simulated spectra using the best-fit models without the Gaussian line, keeping the same responses and background of the observed data. We then fitted each simulated spectrum by adding a Gaussian line and performing a scan over the 3–10 keV energy range using the XSPEC steppar command with steps of keV. This range was chosen to avoid the drop in NuSTAR ’s effective area at 10 keV, which could compromise the reliability of the process (Peca et al., 2021). The resulting best-fit statistic is compared to the statistic from the model without the absorption line. This allows us to calculate the fraction of cases where the exceeds that found in the real spectrum, i.e., the fraction of cases where a fluctuation in the spectrum has a significance equal or higher than that of the absorption line in the real spectrum. This occurred in only 1.4% of the simulations for Model 1 and 5.0% for Model 2. In other words, the probability that the absorption feature is real is 98.6% for Model 1 and 95.0% for Model 2, which can be interpreted as detection significances of approximately 2.5 and 2.0, respectively.
@@ 3.2 August 2024 observation
Based on these results, a follow-up NuSTAR observation was conducted in August 2024. The August 2024 observation has a much lower net count rate, 0.0330.002 . Compared to the 2023 observation, this indicates a decrease in flux of 34% and in signal-to-noise ratio444Computed as , where and are the source and background counts in the source region. from 30 to 24. We performed the same spectral analysis as described for the 2023 observation. The spectrum and confidence contours are shown in Figure 2. The new observation confirmed the absorption feature, as adding a Gaussian line improved the fit by /2 and 11.1/2 for Model 1 and Model 2, respectively, corresponding to significance levels of and . The AIC criterion further supports the presence of the absorption line, yielding and , indicating a “strong” statistical evidence in favor of its inclusion in both models.
@@ 4.1.2 Photoionization modelling
We then employed a more physically motivated approach by modeling the absorption feature using the XSTAR v2.59 photoionization code (Bautista & Kallman, 2001; Kallman & Bautista, 2001). Following Tombesi et al. (2011), we assumed a power-law with , gas temperature T=106K (Nicastro et al., 1999; Bianchi et al., 2005), and a gas density of cm-3. Solar abundances were adopted from Asplund et al. (2009). We simulated three grids555We used the MPI C++ parallelization tool provided by Danehkar et al. (2018) to run XSTAR. with different turbulence velocities, 100, 1000 and 5000 , and a fixed covering factor of 1 (e.g., Pounds et al., 2018). The additional parameters given by these grids are the column density of the absorbing material (), the redshift of the absorber (), and the ionization parameter (). The latter is defined by (erg s-1 cm), where is the ionizing luminosity of the source integrated between 1 and 1000 Ryd (1 Ryd = 13.6 eV), is the gas density, and is the distance between the absorber and the central source (Tarter et al., 1969).
We obtained the best results using the XSTAR grid with a turbulence velocity of 5000 , which provided both constrained parameters and the most improvement to the fit. The results are summarized in Table 2, and our best-fit result for Model 1 is shown in Figure 4. For both models we obtained consistent results, with an ionization parameter and an absorbing column density for the inflowing material of cm-2, with a redshift of , corresponding to an inflow velocity of 0.15. According to the XSTAR modeling, the absorption feature at this redshift is primarily produced by the Fe xxv K transition. As shown in Figure 4, the observed absorption feature is reproduced by the model as a blend of Fe xxv and Fe xxvi K lines, with Fe xxv K providing the dominant contribution. The inclusion of the XSTAR component improves the fit by cstat/d.o.f. = 7.8/3 and 8.3/3 for Models 1 and 2, respectively, corresponding to confidence for three additional free parameters.
@@ 4.1.4 The physical mechanism that drives the inflow
Under these scenarios, the emergence of the absorption feature in the 2023/2024 NuSTAR data, following its absence in the 2010 Suzaku observation, is consistent with the expected episodic nature of such events, which are predicted to vary on timescales of years or less. Additionally, the accompanying flux variations of 20% between the 2010 Suzaku and 2023 NuSTAR observations, and 34% between the 2023 and 2024 NuSTAR observations, further support these interpretations, as such changes are likely to be expected for disk instabilities or fallback processes (e.g., Ghisellini et al., 2004; Nixon et al., 2012).
@@ 5 Summary and conclusions
The absorption feature was consistently detected at 4.8 keV in both 2023 and 2024 NuSTAR observations. Although the second observation yielded a slightly different best-fit energy (4.5 keV), simulations demonstrate this shift is not statistically significant, reinforcing the consistency of the detections. Individually, each detection has a significance that ranges between 2 and 3, as estimated through MC simulations and other statistical tests. When the two observations are considered together, the significance exceeds 4.
The feature was first fitted with a Gaussian absorption line, combined with the two different continuum models that describe the rest of the spectrum. Based on the best fits, the line is tentatively interpreted as a redshifted Fe xxv K or Fe xxvi K transition, corresponding to an inflow velocity of 0.15–0.20.
@@
When fitting the data with XSTAR photoionization models, the best-fit was obtained using a grid with a turbulence velocity of 5000 , where the absorption feature is dominated by a strong Fe xxv K line, blended with a minor contribution from Fe xxvi. The derived parameters, , cm-2, and an inflow velocity of 0.15, are overall in agreement with previous UFIs studies.
@@
Although our results favor an inflow interpretation, we acknowledge the limitations of the current data. While NuSTAR enabled the detection of the absorption feature, its 400 eV energy resolution prevents distinguishing between closely spaced lines such as Fe xxv and Fe xxvi, and limits our ability to characterize the absorber in greater detail. To address this, we have secured observing time with XMM-Newton (AO 24, PI A. Peca), which will provide significantly improved sensitivity and spectral resolution. Additionally, the new XRISM (XRISM Science Team, 2020) mission will be instrumental in further probing the nature and variability of this feature, and more broadly, in understanding the role of UFIs in AGN accretion and feedback.

### v2

@@ Abstract
We report the serendipitous discovery of an absorption feature at 4.8 keV in the NuSTAR spectra of ESP 39607, a Seyfert 2 galaxy at , observed in May 2023 and August 2024. The feature is detected in both observations with individual significance levels between 2 and 3, computed with multiple statistical methods. The combined probability of detecting it in both observations is 4. The absorption feature is consistent with an ultra-fast inflow (UFI) potentially associated with Fe xxv or Fe xxvi K transitions. The inferred inflow velocity is 0.15–0.20, with an estimated launching radius of 22–89 , depending on the assumed iron transition and whether radiation pressure is accounted for. Photoionization modeling associates the UFI primarily with Fe xxv K absorption, blended with a minor contribution from Fe xxvi K. Alternative explanations, including associations with the warm-hot intergalactic medium or outflows of lighter elements, are investigated but found unlikely. If confirmed, this detection represents a rare example of a UFI, providing valuable evidence into extreme and/or non-standard accretion processes near supermassive black holes. Follow-up observations with higher-resolution X-ray spectroscopy, such as with XMM-Newton or XRISM, will be essential to confirm the nature of this feature and better constrain the physical mechanisms driving it.
@@ 2 Data
In addition to the NuSTAR data analyzed in this work, a 60.7 ks Suzaku/XIS observation from December 2010 is also available for ESP 39607. We used the spectra extracted by Ricci et al. (2017) and refer the reader to that work for details on data reduction. The same dataset is also included and shown in the broad-band spectral fitting analysis presented in Peca et al. (submitted), and we refer to both studies for additional modeling details. Applying the same procedure described in Section 3, we did not find evidence of the absorption feature seen in the NuSTAR spectra within the Suzaku data. The implications of this observation are discussed in Section 4. A summary of the observations analyzed in this work is provided in Table 1.
Telescope | Obs. ID | Date | Exp. | Flux |
NuSTAR | 60860001002 | 2023-05-08 | 21.2 | 1.55 |
NuSTAR | 60161025002 | 2024-08-22 | 21.6 | 1.01 |
Suzaku | 705048010 | 2010-12-19 | 60.7 | 1.89 |
@@ 3.1 May 2023 observation
We started by fitting the 2023 observation, which has a net (i.e., background subtracted) count rate of 0.0500.002 . First, we modeled the source spectrum using a simple model composed of an absorbed power-law with an energy cut-off fixed at 200 keV (e.g., Ricci et al., 2017; Zappacosta et al., 2018) plus a pexrav (Magdziarz & Zdziarski, 1995) component to account for reflected emission. For simplicity, we assumed pure reflection by fixing (e.g., Piconcelli et al., 2011; Zhao et al., 2019; Serafinelli et al., 2024); the negative value directs the model to operate in reflection-only mode (e.g., Alexander et al., 2011; Boorman et al., 2025), while a value of one corresponds to reflection from an infinite slab illuminated by isotropic coronal emission (e.g., Marchesi et al., 2016; Zappacosta et al., 2018). The inclination angle was set to the default value of 60∘ (e.g., Ueda et al., 2007; Zhao et al., 2020), and the photon index was linked to that of the primary power-law component. After verifying consistency between the normalizations of the primary power-law and reflection components, we linked them to the same value to minimize degeneracies (e.g., Liu et al., 2017; Peca et al., 2023b). We refer to this configuration as Model 1. Second, we fitted the source spectrum with the more complex, but physically motivated and self-consistent model, UXCLUMPY (Buchner et al., 2019). We fixed the TORsigma and CTKcover parameters to their standard values, 28 and 0.4, respectively, and the inclination angle to 60∘, as for the first model. We refer to this configuration as Model 2. Results from Model 2 are consistent with those from Model 1 within the uncertainties (see Table 2), and both models confirm the obscured and luminous nature of the source, with derived cm-2 and erg s-1, in agreement with those previously reported by Ricci et al. (2017).
The fit residuals revealed a possible absorption feature at 4.8 keV (Figure 1, left panel). To model this feature, we introduced a Gaussian line component to both models, where line normalization and energy were left free to vary. The line width was fixed to eV (e.g., Tombesi et al., 2010; Vignali et al., 2015), as it was found to be unconstrained at the lower end when left free, i.e., consistent with a narrow and unresolved feature. We used the XSPEC steppar command to map the confidence contours on line normalization and energy, and show them in Figure 2. The inclusion of this absorption line improved the fit significantly. For Model 1, we obtained /2, which corresponds roughly to a fit improvement of 3 assuming Gaussian approximation (e.g., Tozzi et al., 2006). For Model 2, the improvement was /2, i.e., 2.7. Model 1 yielded keV and an equivalent width of keV, with Model 2 giving consistent results. We summarize all the detailed results for all model configurations in Table 2.
@@
To further assess the reliability of the absorption feature detection, we adopted a Monte Carlo (MC) approach, as widely used in the literature to evaluate the significance of emission and absorption lines (e.g., Markowitz et al., 2006; Vignali et al., 2015; Costanzo et al., 2022; Peca et al., 2023a). We simulated spectra using the best-fit models without the Gaussian line, keeping the same responses and background of the observed data. We then fitted each simulated spectrum by adding a Gaussian line and performing a scan over the 3–10 keV energy range using the steppar command with steps of keV. This range was chosen to avoid the drop in NuSTAR ’s effective area at 10 keV, which could compromise the reliability of the process (Peca et al., 2021). The resulting best-fit statistic is compared to the statistic from the model without the absorption line. This allows us to calculate the fraction of cases where the exceeds that found in the real spectrum, i.e., the fraction of cases where a fluctuation in the spectrum has a significance equal or higher than that of the absorption line in the real spectrum. This occurred in only 1.4% of the simulations for Model 1 and 5.0% for Model 2. In other words, the probability that the absorption feature is real is 98.6% for Model 1 and 95.0% for Model 2, which can be interpreted as detection significances of approximately 2.5 and 2.0, respectively.
@@ 3.2 August 2024 observation
Based on these results, a follow-up NuSTAR observation was conducted in August 2024. The August 2024 observation has a much lower net count rate, 0.0330.002 . Compared to the 2023 observation, this indicates a decrease in flux of 35% and in signal-to-noise ratio444Computed as , where and are the source and background counts in the source region. from 30 to 24. We performed the same spectral analysis as described for the 2023 observation. The spectrum and confidence contours are shown in Figure 1 (right panel) and Figure 2, respectively. The new observation confirmed the absorption feature, as adding a Gaussian line improved the fit by /2 and 11.1/2 for Model 1 and Model 2, respectively, corresponding to significance levels of and . The AIC criterion further supports the presence of the absorption line, yielding and , indicating a “strong” statistical evidence in favor of its inclusion in both models.
@@ 4.1.2 Photoionization modelling
We then employed a more physically motivated approach by modeling the absorption feature using the XSTAR v2.59 photoionization code (Bautista & Kallman, 2001; Kallman & Bautista, 2001). Following Tombesi et al. (2011), we assumed a power-law with , gas temperature T=106K (Nicastro et al., 1999; Bianchi et al., 2005), and a gas density of cm-3. Solar abundances were adopted from Asplund et al. (2009). We simulated three grids555We used the MPI C++ parallelization tool provided by Danehkar et al. (2018) to run XSTAR. with different turbulence velocities, 100, 1000 and 5000 , and a fixed covering factor of 1 (e.g., Pounds et al., 2018). The additional free parameters given by these grids are the column density of the absorbing material (), the redshift of the absorber (), and the ionization parameter (). The latter is defined by (erg s-1 cm), where is the ionizing luminosity of the source integrated between 1 and 1000 Ryd (1 Ryd = 13.6 eV), is the gas density, and is the distance between the absorber and the central source (Tarter et al., 1969).
We obtained the best results using the XSTAR grid with a turbulence velocity of 5000 , which provided both constrained parameters and the most improvement to the fit. The results are summarized in Table 3, and our best-fit result for Model 1 is shown in Figure 4. For both models we obtained consistent results, with an ionization parameter and an absorbing column density for the inflowing material of cm-2, with a redshift of , corresponding to an inflow velocity of 0.15. According to the XSTAR modeling, the absorption feature at this redshift is primarily produced by the Fe xxv K transition. As shown in Figure 4, the observed absorption feature is reproduced by the model as a blend of Fe xxv and Fe xxvi K lines, with Fe xxv K providing the dominant contribution. The inclusion of the XSTAR component improves the fit by cstat/d.o.f. = 7.8/3 and 8.3/3 for Models 1 and 2, respectively, corresponding to confidence for three additional free parameters. The slightly lower significance, compared to the models with the Gaussian line in absorption, arises because the XSTAR grid introduces three free parameters (, , and ) versus only two for the Gaussian line (line energy and normalization). Additionally, XSTAR accounts for all resonance absorption lines and edges from elements up to in multiple ionization states, among which only the feature presented in this work is statistically significant in our NuSTAR dataset.
@@ 4.1.4 The physical mechanism that drives the inflow
Under these scenarios, the emergence of the absorption feature in the 2023/2024 NuSTAR data, following its absence in the 2010 Suzaku observation, is consistent with the expected episodic nature of such events, which are predicted to vary on timescales of years or less. Additionally, the accompanying flux variations of 20% between the 2010 Suzaku and 2023 NuSTAR observations, and 35% between the 2023 and 2024 NuSTAR observations, further support these interpretations, as such changes are likely to be expected for disk instabilities or fallback processes (e.g., Ghisellini et al., 2004; Nixon et al., 2012).
@@ 5 Summary and conclusions
The absorption feature was consistently detected at 4.8 keV in both 2023 and 2024 NuSTAR observations. Although the second observation yielded a slightly different best-fit energy (4.5 keV), simulations demonstrate that this shift is not statistically significant, reinforcing the consistency of the detections. Individually, each detection has a significance that ranges between 2 and 3, as estimated through MC simulations and other statistical tests. When the two observations are considered together, the significance exceeds 4.
The absorption feature was first fitted with a Gaussian absorption line, combined with two different continuum models that describe the rest of the spectrum. Both models yielded consistent results. Based on the best fits, the line is tentatively interpreted as a redshifted Fe xxv K or Fe xxvi K transition, corresponding to an inflow velocity of 0.15–0.20.
@@
When fitting the data with XSTAR photoionization models, the best-fit was obtained using a grid with a turbulence velocity of 5000 , where the absorption feature is dominated by a strong Fe xxv K line, blended with a minor contribution from Fe xxviK. The derived parameters, , cm-2, and an inflow velocity of 0.15, are overall in agreement with previous UFIs studies.
@@
Although our results favor an inflow interpretation, we acknowledge the limitations of the current data. While NuSTAR enabled the detection of the absorption feature, its 400 eV energy resolution prevents distinguishing between closely spaced lines such as Fe xxv and Fe xxvi K transitions, and limits our ability to characterize the absorber in greater detail. To address this limitation, we have secured observing time with XMM-Newton (AO 24, PI A. Peca). This XMM-Newton observation is expected to provide significantly improved sensitivity and spectral resolution ( eV), making it ideally suited for confirming the absorption feature and better constraining the blended Fe xxv and Fe xxvi. Furthermore, the newly launched XRISM mission (XRISM Science Team, 2020), with its unprecedented energy resolution (5 eV), could enable detailed characterization of the absorber. This would be especially valuable if the intrinsic turbulent velocity is lower than measured here, potentially allowing XRISM to resolve the individual Fe xxv and Fe xxvi components. Together, these missions will offer complementary insights into the nature and variability of this feature, thereby improving our broader understanding of the role that UFIs play in AGN accretion and feedback processes.
acknowledgments
We acknowledge the anonymous referee for the valuable comments that improved the quality of the paper. A.P. acknowledges D. Costanzo for the useful discussions. A.P. and M.K. thank the NuSTAR team for the helpful discussions on the observations and the data reduction of ESP 39607. A.P. and M.K. acknowledge support from NASA through ADAP award 80NSSC22K1126 and NuSTAR grants 80NSSC22K1933 and 80NSSC22K1934. R.S. acknowledges funding from the CAS-ANID grant number CAS220016. C.R. acknowledges support from Fondecyt Regular grant 1230345, ANID BASAL project FB210003, and the China-Chile joint research fund.